# The Eye as a Window to Systemic Health: A Survey of Retinal Imaging from Classical Techniques to Oculomics

### v1

@@ III-C Automated Retinal Vascular Morphology: A Pathway to Oculomics
The circulatory system plays a critical role in systemic health, and the dysfunction often manifests in the retina’s microvascular structures. Retinal circulation, characterised by the unique accessibility and transparency, provides an invaluable opportunity to study vascular health non-invasively. Blood viscosity, vessel length, and vessel diameter are critical determinants of vascular resistance and have been closely examined within the retinal vasculature to uncover systemic disease mechanisms[120]. In connection with these features and parameters, a vascular structure encased in the retinal circulation gained close attention to unveil the mechanism underlying these structures. In recent decades, interest has gone beyond the classical techniques of oculomics due to its relationship with systemic diseases like both types of diabetes, CVDs and dementia[121, 122].
@@
Microvasculature features of the eye in the vivo assessment of FP allow the computerised approach to measuring the retinal vessel calibre of the arterioles, venules, and the ratio between these. These measurements are the gold standard for oculomics, which can link with systemic diseases[50]. Studies showed that hypertension and atherosclerosis are linked with the narrowing of the arteries[120],[123], and the DR is associated with the dilation of retinal veins[124]. In addition, increasing touristy of the retinal arteries is interlinked with hypercholesterolemia and hypertension[125, 126].
@@
Initially, retinal calibres were measured manually, a tedious process requiring human computation and traditional formulas. However, manual vessel segmentation often produces suboptimal results, which can hinder model performance. Over the past few decades, extensive research has been conducted to automate this process and extract retinal vascular features. Significant advancements have been achieved through various approaches, including feature-based methods, unsupervised graph-based techniques, and supervised deep learning models[127, 128, 129, 130, 131, 132, 133]. These developments have led to the emergence of multiple software tools for clinical and research applications. However, most of these tools remain semi-automated, necessitating human intervention to correct vessel segmentation and identify features such as bifurcations, tortuosity, branching angles, artery/vein diameters, widths, and FD.
@@
To address these limitations, AutoMorph, a fully automated software package, was developed by Yukun Zhou et al [1], leveraging DL models to autonomously measure retinal calibres without human intervention. The AutoMorph pipeline consists of four key components: preprocessing, image quality grading, anatomical segmentation, and morphological feature measurement. The anatomical segmentation and morphological feature extraction stages play a crucial role in advancing oculomics, an emerging field that explores the relationship between retinal and systemic diseases. The third stage of AutoMorph enables precise segmentation of the optic disc/cup, vessels, arteries, and veins. The final stage extracts key vascular morphology metrics, including tortuosity (via distance measurement, squared curvature, and density) for each artery and vein, and fractal dimension to quantify vessel complexity. Additionally, it measures retinal vascular calibres for central retinal arterioles equivalent (CRAE) and central retinal venular equivalent (CRVE), along with their arteriolar-to-venular ratio (AVR), and enables macular region-specific assessments.
@@
With advancements in retinal imaging software, researchers can now efficiently quantify retinal vascular features, segment major arteries and veins, and capture the fine-grained structural details of smaller vessels with high precision. These developments have expanded the potential of retinal imaging beyond ophthalmic disease diagnosis, providing valuable insights into systemic health conditions. Increasing evidence suggests that retinal vascular morphology serves as a biomarker for various systemic diseases, including hypertension, diabetes, cardiovascular disease, and neurodegenerative disorders. By enabling automated, high-throughput retinal vascular analysis, AutoMorph and similar technologies contribute to a deeper understanding of the intricate connections between retinal changes and systemic pathophysiology, reinforcing the growing role of oculomics in predictive medicine and disease monitoring. page
Figure 5: Classical Techniques in Eye Research : Classical diagnostic workflow in early eye research, focusing on image preprocessing and handcrafted feature analysis. This figure establishes the foundation for how early computational methods approached retinal diagnostics before oculomics.
Figure 6: Current Techniques Using AutoMorph: Contemporary oculomics framework leveraging AI-based vascular feature extraction tools like AutoMorph. It illustrates the shift toward automated and quantifiable retinal phenotyping to support both ocular and systemic disease research.
Figure 7: Future Prospects in Oculomics: Future outlook for oculomics as a multi-modal healthcare platform integrating imaging, genomics, and computational biology. The figure outlines a vision for system-level precision diagnostics driven by AI and omics technologies.
@@ III-D Evolution of Oculomics: From Classical to Future Frameworks
Figure 7 demonstrates the current framework in oculomics, pitched by automated DL-based image segmentation tools such as AutoMorph. This phase is portrayed by automated vascular feature extraction from retinal images, allowing researchers to quantify high-resolution phenotypic traits. AutoMorph provides structured morphological outputs and objective measures such as CRAE, CRVE, AVR, tortuosity, and fractal dimension. These features not only support ocular disease classification, such as for glaucoma or diabetic retinopathy, but also serve as non-invasive alternatives for systemic conditions like hypertension, cardiovascular disease, and neurological degeneration. This middle phase in the oculomics evolution enables large-scale population analysis and deep phenotyping at scale.
Figure 7 envisions the future direction of oculomics, where phenotypic data from retinal imaging are integrated with molecular-level information to create a fully multi-modal healthcare platform. This framework combines imaging biomarkers with genomic, transcriptomic, proteomic, and metabolomic data layers. Technologies such as genome-wide association studies, protein-protein interaction networks, CRISPR-based gene editing, and pathway enrichment models are embedded within this vision. This integrated pipeline aims to enable precise risk stratification, identify novel therapeutic targets, and personalise disease prediction and management.
@@ IV-A The Eye as a Mirror of Systemic Health
The eyes are considered a window for the mirror of the soul of health. What does this mean to us? This means that looking at the eyes of the person can be interpreted, and the other person’s concealed attitudes, thoughts, and emotions can be guessed [134]. The same is true; the eyes are the only organ in the human body that acts as a window to see the world, making them unique organs to care for. In connection with other non-communicable diseases like rheumatic, neurodegenerative, hypertensive and diabetic Mellitus, eyes will help us to detect them in a non-invasive manner and act as biomarkers for these fatal diseases.
The World Health Organisation estimated that 2.2 billion people in the world are suffering from eye diseases and are being prevented to the tune of one billion of them. Therefore, other than an ophthalmologist, researchers and scientists are keenly interested in unveiling the hidden patterns and mechanisms to analyse with the latest research, emulsified with the holistic approach to care for our eye health and save the whole body [135]. In addition, a recent study revealed that the eye also manifests a variety of stress diseases in which oxidative stress is crucial. This leads to the onset of different degenerative and chronic disorders such as cancer, atherosclerosis, coronary arteries, kidney, neurological, Respiratory, and Rheumatoid arthritis diseases [96]
The advancement in retinal imaging technology, like the FP, OA, OCT and OCTA, has not only enhanced the diagnostic structure related to sight-threatening retinal diseases but also opened the door to leveraging ocular markers to detect systemic disease. This can further anticipate the onset and act as a non-invasive surrogate biomarker for its grading and treatment response [4]. In a nutshell, the eye is a mirror and window for the soul of health, exhibiting numerous diseases. Here, we review some key diseases from recent articles, particularly focusing on chronic disorders with respect to ageing.
@@ IV-B1 Cardiovascular Diseases
Blood vessels residing in the retinal image can be visualised noninvasively, as they are encased with the features of vessel calibre and offer insight into the systemic health of the cardiovascular system and future risk [6]. The stratifying of CV risk is the key challenge which exacerbates the cause of the mortality rate globally. The identification and management of CVDs risk is calculated by considering various variables such as age, smoking status, body mass index (BMI), cholesterol and glucose level, blood pressure, and gender, induced by the patient’s history and blood samples, leveraging the Framingham score and pooled cohort equations [136]. To strengthen this mechanism further, phenotypic information can mitigate the burden of this disease. In addition, the mentioned features are combined in a manner that identifies the types of risk, which are calculated for 10 years.
The study carried out in the [137] shows that some of the unavailable parameters can exacerbate the function of stratifying the CVDs risk. This study shows that cholesterol values were missing in a 10-year risk score and less than 30% for patients; this lags as fasting blood drawn is needed to obtain these data. In contrast, in some studies, BMI is used as an alternative marker of lipids in the preliminary assessment of CVDs. Thus, retinal imaging may be another surrogate marker to predict cardiovascular risk non-invasively and serve as a diagnostic tool for CVDs, offering insights into systemic vascular health [15].
The retinal vasculature is a microcosm of the body’s circulatory system, and its structure and function changes can indicate underlying cardiovascular conditions. Microvasculature abnormalities like the widening and narrowing of venular and arteriolar are strongly connected with Coronary Artery Disease (CAD), hypertension and atherosclerosis [138]. These changes are frequently detected prior to clinical symptoms of CVDs, making the eye a valuable biomarker for early identification. Hypertensive retinopathy, which is characterised by vascular changes such as arteriolar narrowing, HMs, and EXs, is a direct consequence of systemic hypertension. Studies have shown that people with hypertensive retinopathy face an increased risk of stroke, heart attack, and heart failure [139]. Likewise, the other studies explain that [50] low retinal vascular FD and density are responsible for the high risk of mortality incidence, like renal failure, CVD, hypertension, etc.
Furthermore, a condition associated with hypertension and atherosclerosis called retinal vein occlusion has also been linked with CV events. Studies reveal that patients with retinal vein occlusion had a higher risk of developing CAD and a higher mortality rate compared to the general population [7, 140, 141].

### v2

@@ III-C Automated Retinal Vascular Morphology: A Pathway to Oculomics
The circulatory system plays a critical role in systemic health, and the dysfunction often manifests in the retina’s microvascular structures. Retinal circulation, characterised by the unique accessibility and transparency, provides an invaluable opportunity to study vascular health non-invasively. Blood viscosity, vessel length, and vessel diameter are critical determinants of vascular resistance and have been closely examined within the retinal vasculature to uncover systemic disease mechanisms Nguyen & Wong (2006). In connection with these features and parameters, a vascular structure encased in the retinal circulation gained close attention to unveil the mechanism underlying these structures. In recent decades, interest has gone beyond the classical techniques of oculomics due to its relationship with systemic diseases, namely, both types of diabetes, cardiovascular diseases, and dementia Ebuchi et al. (2022); Brazionis et al. (2023).
Automated analysis of retinal vascular morphology provides objective, quantitative measurements of the retinal circulation. Standardized protocols typically focus on the six largest arterioles and venules emerging from the optic disc, measured within predefined concentric regions, most commonly Zone B (0.5–1.0 disc diameters from the disc margin) and Zone C (1.0–2.0 disc diameters). From these measurements, the central retinal arteriolar equivalent (CRAE) and central retinal venular equivalent (CRVE) are derived, summarizing average vessel caliber, while their ratio (AVR) reflects the arteriovenous balance Zhou et al. (2022). Additional quantitative traits extend this characterization: fractal dimension describes the geometric complexity of the branching network Avakian et al. (2002); tortuosity quantifies the curvature and deviation of vessels from a straight path Grisan et al. (2008); vessel density captures the proportion of retinal area occupied by vasculature Curcio & Kar (2019); branching angle measures bifurcation geometry; and arteriovenous crossing patterns define vessel interactions Ikram et al. (2013). These retinal vascular traits provide reproducible phenotypes that can be computed at scale, forming the foundation for downstream oculomics and multi-omics integration.
@@
The microvasculature features of the eye in the in vivo assessment of imaging modalities allow a computerized approach to measuring the caliber of the retinal vessels of the arterioles, venules and the relationship between these. These measurements are the gold standard for oculomics, which can be correlated with systemic diseasesZekavat et al. (2022). Studies showed that hypertension and atherosclerosis are associated with narrowing of the arteriesNguyen & Wong (2006),Cheung et al. (2007), and DR is associated with dilation of the retinal veinsWong (2011). In addition, increasing retinal artery tortuosity is interrelated with hypercholesterolemia and hypertensionCheung et al. (2011); Owen et al. (2011).
@@
Initially, retinal calibers were measured manually, a tedious process requiring human computation and traditional formulas. However, manual vessel segmentation often produces suboptimal results, which can hinder model performance. Over the past few decades, extensive research has been conducted to automate this process and extract retinal vascular features. Significant advancements have been achieved through various approaches, including feature-based methods, unsupervised graph-based techniques, and supervised deep learning modelsMirsharif et al. (2013); Dashtbozorg et al. (2013); Estrada et al. (2015); Ronneberger et al. (2015); Huang et al. (2018); Srinidhi et al. (2019); Hatamizadeh et al. (2022). These developments have led to the emergence of multiple software tools for clinical and research applications. However, most of these tools remain semi-automated, necessitating human intervention to correct vessel segmentation and identify features such as bifurcations, tortuosity, branching angles, artery/vein diameters, widths, and Fractal dimension.
@@
To address these limitations, AutoMorph, a fully automated software package, was developed by Zhou et al. (2022), leveraging deep learning models to autonomously measure retinal calibers without human intervention. The AutoMorph pipeline consists of four key components: preprocessing, image quality grading, anatomical segmentation, and morphological feature measurement. The anatomical segmentation and morphological feature extraction stages play a crucial role in advancing oculomics, an emerging field that explores the relationship between retinal and systemic diseases. The third stage of AutoMorph enables precise segmentation of the optic disc/cup, vessels, arteries, and veins. The final stage extracts key vascular morphology metrics, including tortuosity (via distance measurement, squared curvature, and density) for each artery and vein, and fractal dimension to quantify vessel complexity. Additionally, it measures retinal vascular calibers for CRAE and CRVE, AVR, and enables macular region-specific assessments.
@@
With advancements in retinal imaging software, researchers can now efficiently quantify retinal vascular features, segment major arteries and veins, and capture the fine-grained structural details of smaller vessels with high precision. These developments have expanded the potential of retinal imaging beyond ophthalmic disease diagnosis, providing valuable insights into systemic health conditions Hanssen et al. (2022). Increasing evidence suggests that retinal vascular morphology serves as a biomarker for various systemic diseases, including hypertension, diabetes, cardiovascular disease, and neurodegenerative disorders Honavar (2022). By enabling automated high-throughput retinal vascular analysis, AutoMorph and similar technologies contribute to a deeper understanding of the intricate connections between retinal changes and systemic pathophysiology, reinforcing the growing role of oculomics in predictive medicine and disease monitoring.
@@ III-D Evolution of Oculomics: From Classical to Future Frameworks
Figure 6 represents the current phase of oculomics, where deep learning automates the entire process of retinal vascular analysis. Starting with quality control, fundus images are either accepted or rejected before progressing through vessel probability mapping, binary segmentation, artery–vein separation, skeletonisation, and disc/cup segmentation. Standard epidemiological zones (B and C) are then overlaid around the optic disc to provide consistent reference frames for vascular measurement. From these outputs, quantitative biomarkers such as CRAE, CRVE, AVR, fractal dimension, vessel density, and multiple tortuosity indices are derived in a standardized way. Unlike the handcrafted approaches of earlier decades, these pipelines generate reproducible traits at scale, making them suitable for population cohorts and cross-study comparisons. By translating images into structured, quantitative endophenotypes, automated analysis has transformed the retina into a measurable biomarker source rather than a descriptive image. This stage is crucial in the evolution of oculomics. It bridges classical imaging, which remained focused on ocular disease, and future frameworks, where retinal traits are integrated with genetics, proteomics, metabolomics, and digital health data. In other words, Figure 6 does not just show technical automation; it marks the point where retinal imaging becomes scalable, reproducible, and ready for systemic integration.
Recent studies further illustrate how oculomics is expanding beyond descriptive pipelines into integrative, mechanistic frameworks. For instance, Zekavat et al. (2022); Lu et al. (2025) demonstrated how high-resolution fundus datasets can be combined with clinical records and Mendelian randomization to show that changes in vascular morphology, such as arteriolar narrowing or reduced fractal dimension, are not merely correlates but causally implicated in major vascular events including stroke and myocardial infarction. This move from association to causation provides evidence that retinal traits can function as mechanistic biomarkers rather than statistical surrogates. Deep phenotyping cohorts, such as the UK Biobank Collins (2012) and the 10K Shilo et al. (2021) study, further strengthen this approach by embedding retinal imaging within multi-layered datasets that include lipidomics, metabolomics, immune profiling, body composition, and many more. Complementing these resources, Reicher et al. (2025) has extended deep phenotyping even further by integrating genetics, transcriptomics, metabolomics, immune and microbiome profiling, lifestyle data, continuous physiological monitoring, and imaging into a prospective cohort of more than 28,000 participants, alongside the development of foundation AI models trained for disease prediction. In another study, analysis from the 10K project has shown how deeply phenotyped cohorts can also reveal sex-specific dynamics of biological ageing, with machine learning models generating system-specific biological age scores that predict age-related disease risk beyond chronological age Reicher et al. (2024). These studies reveal that retinal traits can be contextualised within broader biological signatures of health and ageing, providing a richer framework for interpreting systemic variation across populations.
In parallel, plasma proteomic research has highlighted how organ-specific proteomic ages, particularly those of the brain and immune system, are strong predictors of Alzheimer’s disease and longevity, with hazard ratios exceeding 3.0 even after adjustment for conventional risk factors Oh et al. (2025). Such findings offer a mechanistic model for how retinal imaging could be linked with proteomic and genomic data to stratify risk for neurodegeneration and systemic aging. Disease-specific pipelines also illustrate the breadth of this vision: Eye2Gene, trained on multimodal retinal imaging, achieved top-five diagnostic accuracy of nearly 84% across 63 inherited retinal disease genes, demonstrating how foundation-style phenotyping can accelerate genetic diagnosis Pontikos et al. (2025). Similarly, DeepSLE, a deep learning system to detect Systemic lupus erythematosus (SLE), has shown that autoimmune disease, specifically SLE, can be detected from fundus photographs using AI with performance comparable to serological screening, extending oculomics beyond cardiometabolic and neurodegenerative domains into immunology Li et al. (2025).
Together, these examples capture the trajectory summarised in Figure 7. Current studies are beginning to show how causal inference can elevate retinal biomarkers from correlates to mechanisms, how multimodal concatenation of retinal and clinical data improves prediction, and how foundation-style models like Eye2Gene can compress complex phenotypes into actionable diagnostic outputs. At the same time, multi-system screening models such as DeepSLE preview a future where a single retinal image could support simultaneous risk profiling across cardiovascular, neurological, and immune conditions. Looking forward, the integration of retinal imaging with genomic-wide associations, plasma proteomics, metabolomics, digital devices, and longitudinal clinical data, processed through ensemble and foundation AI models, will define the next-generation framework of oculomics. In this vision, the inputs from genomics and clinical/digital streams (Figure 7A and 7C) converge in advanced computational architectures (Figure 7B) to produce validated biomarkers, risk stratification tools, and decision-support systems that move oculomics firmly into the domain of precision medicine.
@@
Figure 5: Before the rise of oculomics, computational pipelines for ophthalmology followed a straightforward path: Images were acquired as input for clarity, processed through pre-processing. The classical method went through feature engineering with a need for manual interpretation, including contrast enhancement, region of interest cropping and segmentation, and noise reduction. Then these were put in the model for training, testing and validating the model for the specific ocular diseases for machine learning and early deep learning modelling. This work is in the classification, detection and prediction of different diseases. These systems focused almost entirely on ocular diseases such as cataract, glaucoma, age-related macular degeneration, diabetic retinopathy, and corneal opacity. While they laid the foundation for computer-assisted diagnosis, they were limited in scope, largely handcrafted, and confined to the eye itself without considering wider systemic health.
Figure 6: (A) Left-to-right workflow: Fundus images undergo automated quality assessment, followed by vessel segmentation (binary masking), artery/vein classification, skeletonization, and disc/cup delineation. Peripapillary measurement zones (B & C) are defined as concentric annuli centred on the optic disc for standardized caliber and trait quantification. (B) Vascular Biomarkers: The system extracts quantitative vascular metrics, including dimensions, fractal dimension, vessel density, and tortuosity indices for both arteries and veins. Outputs, generated using AutoMorph on CHASE DB1 (image 007-L) Owen et al. (2009), are reported in micrometres calibrated to the device’s field of view. (C) Transition to Systemic Biomarkers: By standardizing and scaling retinal phenotypes, AI facilitates the association of ocular conditions (e.g., diabetic retinopathy, glaucoma, age-related macular degeneration) with systemic diseases (cardiovascular, metabolic, neurodegenerative), transitioning from handcrafted descriptors to traits suitable for population studies and enabling multi-omics integration.
Figure 7: Future prospects in oculomics, toward validated biomarkers and precision medicine. (A) Genomic & molecular inputs: GWAS, gene expression, pathways, protein–protein interactions, single-nucleotide polymorphisms (SNPs), and gene-editing constraints provide a mechanistic context for retinal traits. (C) Clinical, environmental & digital context: metabolomics, microbiome, medical history, lifestyle and environmental factors, histo-molecular readouts, and data from digital devices. (B) Computational core: multi-modal data fusion supports four complementary strategies, (i) causal inference (e.g., Mendelian randomization) to separate correlation from causation; (ii) multimodal ensembles/concatenation for robust prediction; (iii) foundation/next-generation phenotyping models (e.g., Eye2Gene for IRDs) to learn generalizable retinal representations; and (iv) multi-system screening models that jointly profile ocular and systemic phenotypes. Outputs include validated biomarkers, risk stratification, and decision support for precision medicine. Example systems span automated vascular phenotyping (AutoMorph), phenotypic screening with genetic validation for vascular events, deep learning for systemic autoimmunity (DeepSLE), and IRD gene prioritization (Eye2Gene).
@@ IV-A The Eye as a Mirror of Systemic Health
The eyes are considered a window for the mirror of the soul of health. What does this mean to us? This means that looking at the eyes of the person can be interpreted, and the other person’s concealed attitudes, thoughts, and emotions can be guessed Ph.D (2024). The same is true; the eyes are the only organ in the human body that acts as a window to see the world, making them unique organs to care for. In connection with other non-communicable diseases like rheumatic, neurodegenerative, hypertensive and diabetic Mellitus, eyes will help us to detect them in a non-invasive manner and act as biomarkers for these fatal diseases.
The World Health Organisation estimated that 2.2 billion people in the world are suffering from eye diseases and are being prevented to the tune of one billion of them. Therefore, other than an ophthalmologist, researchers and scientists are keenly interested in unveiling the hidden patterns and mechanisms to analyse with the latest research, emulsified with the holistic approach to care for our eye health and save the whole body XLab Health (2024). In addition, a recent study revealed that the eye also manifests a variety of stress diseases in which oxidative stress is crucial. This leads to the onset of different degenerative and chronic disorders such as cancer, atherosclerosis, coronary arteries, kidney, neurological, Respiratory, and Rheumatoid arthritis diseases Honavar (2022); Wei et al. (2022)
The advancement in retinal imaging technologies, such as fundus photography, OCT, and OCTA, has not only improved the diagnosis of sight-threatening ocular diseases but also opened the door to using ocular markers as indicators of systemic conditions. These biomarkers can anticipate disease onset and serve as non-invasive surrogates for risk stratification and treatment monitoring Wagner et al. (2020). In essence, the eye is both a mirror and a window to systemic health, reflecting numerous chronic disorders. Building on this perspective, the following section reviews systemic diseases where retinal changes are most evident, with a particular emphasis on age-related conditions.
@@ IV-B1 Cardiovascular Diseases
Blood vessels captured in retinal images can be visualised noninvasively, and subtle differences in their caliber and branching patterns provide valuable insights into systemic vascular health and future cardiovascular risk Hanssen et al. (2022). The ability to stratify risk is central to preventing CVD, which remains a leading cause of mortality worldwide. Current clinical scores, such as the Framingham Risk Score (FRS) and pooled cohort equations, rely on factors like age, smoking, body mass index (BMI), cholesterol, glucose, blood pressure, and sex, often supplemented by blood tests Goff Jr et al. (2014). However, these approaches face practical limitations. For instance, cholesterol values are not always available, with one study noting they were missing for nearly 30% of patients in a 10-year risk calculation Hira et al. (2015). Alternatives, such as using BMI in place of lipids, have been proposed, but they only partially capture metabolic risk. In this context, the retina offers an attractive non-invasive surrogate, providing a direct view of the microvasculature and its relationship to systemic disease Poplin et al. (2018).
Early population studies consistently reported that narrower arterioles and wider venules were associated with hypertension, coronary artery disease, and atherosclerosis Cheung et al. (2012). These cross-sectional observations highlighted the eye’s potential as a cardiovascular biomarker but did not establish causality. More definitive evidence came from large-scale longitudinal analyses. A pooled individual-participant meta-analysis of 20,798 people across six cohorts showed that wider venular caliber independently predicted incident stroke, with a hazard ratio of 1.15 (95% CI: 1.05–1.25) per $20\,\mu\text{m}$ increase, whereas arteriolar caliber was not significantly associated (HR 1.00, 95% CI: 0.92–1.08) McGeechan et al. (2009). Importantly, incorporating venular caliber into standard risk models reassigned more than 10% of individuals at intermediate stroke risk, underscoring the added predictive value of retinal vascular traits. Similarly, hypertensive retinopathy,characterised by vessel narrowing, haemorrhages, and exudates, not only reflects chronic blood pressure elevation but also signals a higher likelihood of stroke, heart attack, and heart failure Wang, Feng, Wu, Ding, Liu, Xie & Ma (2024). These findings underscore that the vascular features of the retina can serve as validated predictors, not merely correlates, of cardiovascular outcomes.
More recently, advances in artificial intelligence have expanded the predictive scope of retinal imaging. A deep learning model was induced by Rudnicka et al. (2022) leveraging retinal vascular traits to predict major adverse cardiovascular events, including circulatory mortality, myocardial infarction, and stroke, with a C-statistic of 0.75 to 0.77, comparable to FRS. In a similar study, Poplin et al. (2018) showed that fundus photographs alone could predict 5-year cardiovascular events with an AUC of 0.70, matching the performance of the composite SCORE calculator. In people with diabetes, the addition of retinal parameters to polygenic risk scores for coronary artery disease improved prognostic accuracy over a 10-year period Mordi et al. (2022). Other biomarkers, such as the retinal age gap, a measure comparing biological and chronological age derived from retinal images, have also been shown to predict cardiovascular mortality in multiple cohorts Nusinovici et al. (2022); Zhu et al. (2023).
Integrative approaches further highlight the potential of oculomics. Combining fundus photographs with conventional clinical risk factors or dual-energy X-ray absorptiometry improved CVD risk prediction by 2–3% Al-Absi et al. (2022); Mellor et al. (2023). Optical coherence tomography angiography-based vascular parameters have also been investigated as markers of coronary artery disease Zhong et al. (2022). Although their additional predictive value over established risk factors remains modest, particularly in patients with advanced diabetic or hypertensive retinopathy, they demonstrate the breadth of retinal imaging modalities that can contribute to cardiovascular risk assessment.
Finally, large-scale AI studies integrating genetics have begun to provide causal insights Zekavat et al. (2022). In the UK Biobank, deep learning models predicted 10-year risks of stroke, myocardial infarction, and chronic kidney disease with AUC values (area under the receiver operating characteristic curve) of 0.74–0.77, performing on par with or better than traditional clinical scores. Importantly, genome-wide association studies and Mendelian randomization analyses confirmed causal links between retinal vascular traits and cardiovascular outcomes, while also identifying reverse effects of chronic kidney disease on retinal microcirculation Lu et al. (2025). These advances mark a shift from descriptive correlations to validated, mechanistic pathways, placing the retina at the forefront of precision cardiovascular medicine.